OpenStudy (anonymous):

if sin theta= 4/5 and theta is in quadrant 2 what is cot theta

OpenStudy (anonymous):

sin x=4/5=p/h so h=5 and p=4 tan x = p/b using Pythagoras Theorem h²=p²+b² b=√(h²-p²)=√(5²-4²)=√(25-16)=√9=3 tan x =p/b=4/3 Because tan is -ve in second quadrant tan x = -4/3 answer

OpenStudy (mertsj):